The sum of two numbers is 37. The difference of these same numbers is 11. Find the numbers. Must show work
Answer by scott8148(6628)

x + y = 37

x - y = 11

adding equations together (eliminating y) ___ 2x = 48 ___ x = 24

substituting ___ (24) + y = 37 ___ y = 13

check ___ (24) - (13) = 11